# Robust Identity Perceptual Watermark Against Deepfake Face Swapping

### v2

@@ 4.6 Ablation Study on Common Manipulations
In this paper, following the convention of steganography, we picked the common manipulations that best imitate the unavoidable image quality degradation of posting and spreading in real-life scenarios and set the parameter values with considerable difficulties. Particularly, the parameters are set equal to or harder than the early work [61, 18, 28] in our experiments. In this section, to further validate the robustness of the proposed watermarking framework, we conducted ablation testing sessions on CelebA-HQ at the $256\times 256$ resolution with easier and harder parameters that are unseen during training for the common manipulations, respectively. As a result, in addition to the outstanding performance reported in Sections 4.2 and 4.3, Table 6 implies that the robustness is completely guaranteed for common manipulations with easier parameter values, and it is also only slightly affected in harder conditions, with all accuracies still above $90\%$. In conclusion, the potential effects of common manipulations at different difficulty levels are trivial with respect to the robustness of our proposed approach.
5 Conclusion
@@
In this paper, we propose a robust identity perceptual watermarking framework that proactively defends against malicious Deepfake face swapping. By securely assigning identity semantics to the watermarks, our approach concurrently accomplishes detection and source tracing with a single robust watermark for the first time. Meanwhile, the confidentiality is guaranteed by devising an unpredictable chaotic encryption system on the identity perceptual watermarks. Extensive experimental results have proved that our method is robust when facing both common and malicious manipulations on the watermarked images even under cross-dataset settings. Moreover, by leveraging a single Deepfake face swapping algorithm for fine-tuning, we are able to preserve promising cross-manipulation performance against unseen face swapping algorithms. The AUC scores of 98.66% and 98.97% at $128\times 128$ and $256\times 256$ resolutions further prove the promising detection ability of our approach on Deepfake face swapping, significantly outperforming the state-of-the-art passive detectors.

### v3

@@ IV-G Ablation Studies
Common Manipulations at Various Difficulties. In this paper, following the convention of steganography, we picked the common manipulations that best imitate the unavoidable image quality degradation of posting and spreading in real-life scenarios and set the parameter values with considerable difficulties. Particularly, the parameters are set equal to or harder than the early work [34, 35, 36] in our experiments. In this section, to further validate the robustness of the proposed watermarking framework, we conducted ablation testing sessions on CelebA-HQ at the $256\times 256$ resolution with easier and harder parameters that are unseen during training for the common manipulations, respectively. As a result, in addition to the outstanding performance reported in Sections IV-B and IV-C, Table IX implies that the robustness is completely guaranteed for common manipulations with easier parameter values, and it is also only slightly affected in harder conditions, with all accuracies still above $90\%$. In conclusion, the potential effects of common manipulations at different difficulty levels are trivial with respect to the robustness of our proposed approach.
Watermark Confidentiality Analyses. A chaotic encryption system is devised to securely protect the watermarks in Section III-B to avoid potential adaptive attacks. In particular, an attacker may attempt to reversely deduce the original watermark based on an encrypted one. In our approach, following the parameter settings in Section IV-A, the encryption pipeline becomes unpredictable and nonreversible relying on Eqn.(2). In Table X, besides fixing $x_{0}$ at the chosen value $0.1$ in this study, we listed the sequence of $x_{i}$ for $1<=i<=30$ by attempting different $r$ values. It can be observed that, after a few iterations, $x_{i}$ starts to oscillate between cyclic values for $r=2$, $r=2.5$, and $r=3.5$. Specifically, $x_{i}$ stables at $0.5$ (red) for $r=2$ and varies around $0.59$ (red) and $0.60$ (blue) for $r=2.5$. As for $r=3.5$, four oscillation patterns can be concluded between approximately $0.86$ (red), $0.40$ (blue), $0.84$ (green), and $0.46$ (brown). On the contrary, the evolution of $x_{i}$ is completely unpredictable with no patterns observed for the chosen $r=3.93$ in this study. In other words, with no evolution pattern in the chaotic sequence, the adaptive attacks are prevented.
Figure 4: Bifurcation diagram of the logistic map Eqn.(2) for continuous $r$ values with $x_{0}=0.1$.
Figure 5: Bifurcation diagrams with different $x_{0}$ values.
In Fig. 4, we visualized the bifurcation diagram of the logistic map of Eqn.(2) with $x_{0}=0.1$ for a continuous range of $0<=r<=4$. In particular, by omitting $x_{i}$ for $0<=i<10$, the $200$ consecutive $x_{i}$ values are plotted as points for each $r$ value after deriving $150$ linear spaces within the range. Moreover, we zoomed in on the view to $3<=r<=4$ for a closer observation of the oscillation and unpredictable behaviors of $x_{i}$ at different $r$ values. It can be concluded that the chaotic characteristic in $X$ occurs for approximately $r>3.6$ after periodic sequences for lower $r$ values.
Besides, we presented the visualizations with different $x_{0}$ values in Fig. 5, which proves that the choice of $x_{0}$ does not affect the chaotic behavior. Therefore, selecting a proper $r$ value with a chaotic sequence $X$ for the generation of the encryption key following Eqn.(3) makes the watermark encryption pipeline unpredictable and nonreversible even though the adversarial attackers may have access to the encrypted watermarks. Meanwhile, it is not necessary to adopt the first $l$ values in $X$ for watermark encryption, because randomly selecting an $x_{i}$ for $i\neq 0$ as the starting point to derive the encryption key $K$ can further ensure watermark safety.
Furthermore, we provided a demonstration of encrypting a sample binary watermark $m_{c}$ of length 20 using the chaotic encryption system following Eqn.(2) and Eqn.(3). To begin with, we defined the sample watermark as follows.
$m_{c}$: 0 1 0 1 0 1 0 0 0 1 1 1 0 0 0 1 0 1 0 0
To encrypt $m_{c}$, we adopted $x_{10}$ through $x_{19}$ with $r=3.93$ from Table X to generate the encryption key sequence $K$. Following Eqn.(3) with constant values $p=5$ and $q=11$, the $x_{i}$ values are transformed to $K$ as follows.
$K$: 0 1 0 0 1 1 1 0 0 0 1 1 1 1 1 1 0 1 0 0
An XOR operation between $m_{c}$ and $K$ derives $m_{e}$ as follows.
$m_{e}$: 0 0 0 1 1 0 1 0 0 1 0 0 1 1 1 0 0 0 0 0
This encrypted watermark $m_{e}$ is confidentially protected against adversarial attackers and can be used for watermark embedding. Contrarily, an authorized individual can easily decrypt $m_{e}$ to derive $m_{d}$ by an extra XOR operation with the encryption key $K$.
$m_{d}$: 0 1 0 1 0 1 0 0 0 1 1 1 0 0 0 1 0 1 0 0
As a result, one can find it faithfully identical to the original watermark $m_{c}$.
V Conclusion
@@
In this paper, we propose a robust identity perceptual watermarking framework that proactively defends against malicious Deepfake face swapping. By innovatively and securely assigning identity semantics to the watermarks, our approach concurrently accomplishes Deepfake detection and source tracing with a single robust watermark for the first time. Meanwhile, confidentiality is guaranteed by devising a chaotic encryption system that brings unpredictability to identity perceptual watermarks. Extensive experimental results have proved that our method is robust when faced against both common and malicious manipulations, even under cross-dataset settings. Moreover, by leveraging a single face swapping algorithm, SimSwap, for fine-tuning, we are able to preserve promising cross-manipulation performance against unseen face swapping algorithms. The AUC scores of $95.39\%$ and $97.57\%$ at $128\times 128$ and $256\times 256$ resolutions against mixed manipulations have further proved the promising detection ability of our approach against Deepfake face swapping, significantly outperforming the state-of-the-art passive detectors. Lastly, the reliable top-5, top-3, and top-1 source tracing accuracies of 98.10%, 97.72%, and 96.56% favorably demonstrate the ability of our method to trace back to the original victim identity along with proactive Deepfake detection.